OpenStudy (twizttiez):

Which number is irrational? A. 3/8 B. 56/8 to the fourth C. 36 to the fourth over 4 to the fourth D. 0.19 repeating

OpenStudy (twizttiez):

3/8= 0.375 56/8 to the fourth =2.64575131... 36 to the fourth over 4 to the fourth=3 And 0.19 is repeating so we can already exclude D

OpenStudy (twizttiez):

@mathmate

OpenStudy (twizttiez):

\[\sqrt{56/8}=2.64575131... \]

OpenStudy (twizttiez):

I think would be my answer am I correct

OpenStudy (twizttiez):

So I'm not? That's not very helpful sorry

OpenStudy (anonymous):

B and C are both irrational